in what time would a body acquire a velocity of 270m/h if it was started with a velocity of 2cm/s and moved with an acceleration of 1/2 cm/s*s ?find the average velocity of the body for its motion during this period in two ways by using (i) the initial and final velocity of the body and (ii)the initial velocity,the acceleration and time?
1
Expert's answer
2012-07-26T08:26:18-0400
Let Vf and V0 be the final and initial velocities respectfully, a be the acceleration and t the elapsed time. Then

Vf = V0 + at ==> t = (Vf - V0)/a.

So,

t = (270 - 0.02)/0.005 = 53996 s.

find the average velocity of the body for its motion during this period in two ways by using
(i) the initial and final velocity of the body

Va = (Vf - V0)/2 = (270 - 0.02)/2 = 134.99 m/s;

(ii)the initial velocity, the acceleration and time?

Va = (Vf - V0)/2 = Va = (V0 + at - V0)/2 = at/2 = 0.005*53996/2 = 134.99 m/s.